You would like to save annually for buying a car 6 years from today. Suppose the first deposit is made today and the last deposit will be made 5 years from now. Assume the car will cost you $30,000 and your deposits earn you interest at 6% p.a. compounded annually.

a. What is your annual deposit amount?

b. Instead of making annual deposits, you would like to make your deposit monthly and the bank is happy to pay your interest on a monthly basis. What is the APR that would make the bank indifferent to these two way of paying interest?

c. Assuming you make the deposit of $300 at the end of each month and the first deposit will be made one month from today, use the answer to part b calculate how much your deposits would accumulate to 6 years later

d. Assume 6 years later, the car price has gone up to $35,000. You decide to use the accumulated deposit as the down payment and take up a 2-year car loan. What would your monthly payment be if the interest rate is 7.5% and compounded monthly?

(Show the working and formulas for each answers)
Answer #1
Rate Annual interest rate 6%
Nper Number of annual deposit 6
Fv Future Value required $30,000
Type Deposit at the beginning of year 1
PMT Annual Deposit Required $4,057.43 (Using PMT function of excel with Rate=6%,Nper=6,Fv=-30000,Type=1)
b Monthly Deposit
Nper Number of monthly deposit =6*12 72
Pmt Monthly deposit =4057/12= $338.12
Fv Future Value required $30,000
RATE Monthly interest rate required 0.57% (Using RATE function of excel withNper=72,Pmt=-338.12Fv=30000)
APR that will make the bank indifferent 6.84% (0.57%*12)
c If $300 is deposited per month
Nper Number of monthly deposit =6*12 72
Rate Monthly interest rate 0.57%
Pmt Monthly deposit $300
FV Accumulated amount at end of 6 years $26,617.82 (Using FV function of excel with Rate=0.57%,Nper=72,Pmt=-300)
d
Down Payment $26,617.82
Pv Loan amount =35000-26617.82= $8,382.18
Rate Monthly interest rate=(7.5/12)% 0.625%
Nper Number of months of loan repayment 24
PMT Monthly Payment Required $377.19 (Using PMT function of excel with Rate=0.625%,Nper=24,Pv=-8382.18)

=PMT(F4,F5,-F6,1) F8 D Н K L. Annual interest rate 6% 4 Rate Number of annual deposit Nper Future Value required $30,000 Fv DClipboard Font Alignment Number Styles Cells Editing F14 =RATE(F11,-F12,F13) Н Annual interest rate 6% 4 Rate Number of annuafe =FV(F19,F18,-F20) F21 D Н K L. 3 Annual interest rate 6% 4 Rate Number of annual deposit 5 Nper Future Value required $30,fe =PMT(F26,F27,-F25) F28 н K M Deposit at the beginning of year Annual Deposit Required Type $4,057.43 (Using PMT function o